Kasey1210:

Simplify: (2x4y3) • (6x3y2)

Nnesha:

\[\large\rm (2x^4y^3) • (6x^3y^2)\] \[\rm ( 2 \cdot 6) (x^4 \cdot x^3)(y^3 \cdot y^2)\] apply the exponent rule \[x^a \cdot x^b = x^{a+b}\] when multiply base we can add their exponents

Ultrilliam:

@kasey1210

Nnesha:

stop spamming or i'll ban u